### v1

@@ 5.1 Experimental Setup
Data Collection & Model Training. We collect production traces from ${Google}$ that consist of the historical execution log of the data processing framework and the I/O traces from the distributed storage system. These traces contain jobs’ metadata and post-execution measurements, such as TCIO. Section 4.2 explains the features we use. Our training and test dataset each contains one week’s data, which are collected from a contiguous two-week time span.
@@ 5.4 Generalizability [RQ4]
We explore generalizability of our method across users and workloads. In practice, good generalizability is necessary as infrastructure, user behaviors, workloads, etc. change over time. First, we evaluate the generalizability across workloads in Figure 9. We train our category model on clusters $C0/C1/C2/C3$, and evaluate their performance on $C0$. $C3$ is a special cluster that only runs certain workloads that are rare in other clusters. We thus find that our method can adapt to unseen workloads, except for outliers.
@@ 5.5 Model Analysis [RQ5]
Another practical requirement for ML models is their execution latency and explainability. We show the accumulated inference time of 50 jobs in Figure 9(a), where running the model takes about 4 ms per job, fast enough for making online placement decisions. Note that our ML model invocation is currently an unoptimized prototype implemented in Python. Potential efficiency improvements could be achieved through further performance optimization, such as using Yggdrasil’s C++ bindings.
@@
We also find that our end-to-end savings may not benefit from more accurate models in Figure 11. Here, we evaluate the performance of the TCO savings assuming we have perfect models. The “Predicted category” is our approach. The “True category” is a method where we replace the category model’s prediction results with the ground truth, representing 100% prediction accuracy. We think this observation can help us rethink the solution space of ML for Systems. The challenges of ML for systems are not only about the accuracy and learning algorithm but also about how to formulate the learning problem and how to use the ML model.
@@ C.2 Sensitivity Analysis
Adaptive Algorithm Parameters. We include all combinations of hyperparameters where $T_{\textsc{SpilloverTCIO}}\in\{[0.005,$ 0.03],0.03],0.03 ] , [0.01,[0.01,[ 0.01 , 0.15],[0.05,0.25]}0.15],[0.05,0.25]\}0.15 ] , [ 0.05 , 0.25 ] }, look back window time length (seconds) $t_{w}\in\{600,900,1800\}$, and admission decision effective time $t_{l}\in\{600,900,$ 1800}1800\}1800 }. We evaluate the sensitivity of the TCO savings for the same set of workloads in Figure 9. For each parameter combination, we apply the same parameter settings to all the workloads in the group. In Figure 15, the blue area in the figure presents the upper bound and lower bound of TCO savings under different SSD capacities across different hyperparameter combinations. Our solution is not sensitive in terms of hyperparameter selection in the adaptive algorithm.
@@ C.3 Adaptive Category Selection Dynamics
To demonstrate the dynamics of our adaptive algorithm, we present the pattern of category threshold change and spill over percentage in Figure 16. We track the threshold change for 1 week online. Our adaptive category selection algorithm can adjust the category admission threshold to a higher range when SSD quota is limited and allow more category admissions when SSD space is plentiful.
@@ Appendix E Discussion
Our approach can be adopted in other deployments and frameworks. The solution flow of our system is: 1) Setting an optimization objective (TCO savings in our case). 2) Designing a ‘hint’ (workload model output) passed from each workload from the application layer to the caching layer. Our hint is job importance in terms of TCO savings. 3) Picking jobs of different categories based on feedback from system utilization.
This flow remains adaptable to other applications. The key components that vary are the objective function, available model features, and how system utilization is quantified.

### v2

@@ 5.1 Experimental Setup
Data Collection & Model Training. We collect production traces from ${Google}$ that consist of the historical execution log of the data processing framework and the I/O traces from the distributed storage system. These traces contain jobs’ metadata and post-execution measurements, such as TCIO. Section 4.2 explains the features we use. Our training and test dataset each contains one week’s data, which are collected from a contiguous two-week time span. We include both read-heavy and write-heavy jobs, reflecting the diversity of workloads across the hyperscale fleet. Block sizes of these workloads vary from several KiB to MiBs.
@@ 5.4 Generalizability [RQ4]
We explore generalizability of our method across users and workloads. In practice, good generalizability is necessary as infrastructure, user behaviors, workloads, etc. change over time. First, we evaluate the generalizability across workloads (Figure 8). We train one category model for each cluster $C0/C1/C2/C3$, and evaluate their performance on $C0$. $C3$ is a special cluster that only runs certain workloads that are rare in other clusters. We thus find that our method can adapt to unseen workloads, except for outliers.
@@ 5.5 Model Analysis [RQ5]
Another practical requirement for ML models is their execution latency and explainability. We show the accumulated inference time of 50 jobs in Figure 9(a), where inference takes about 4 ms per job (considerably lower compared to the 99ms of the Transformer model in Zhou & Maas (2021)), fast enough for making online placement decisions. Note that our ML model invocation is currently an unoptimized prototype implemented in Python. Potential efficiency improvements could be achieved through further performance optimization, such as using YDF’s C++ binding.
@@
We also find that our end-to-end savings may not benefit from more accurate models in Figure 11. Here, we evaluate the performance of the TCO savings assuming we have perfect models for classification prediction. The “Predicted category” is our approach. The “True category” is a method where we replace the category prediction described in Section 4.2 with the ground truth category, representing 100% prediction accuracy.
This observation highlights an important insight: For a given algorithm, there are diminishing returns from further improving model accuracy beyond a certain point. Future improvements in TCO savings are more likely to come from refining the design of the category itself—specifically, its ability to accurately capture job importance—and enhancing its integration with the underlying adaptive algorithm. The synergy between effective category design and the adaptive algorithm plays a crucial role in achieving optimal results.
This finding helps us rethink that the challenges of ML for Systems are not solely about improving model accuracy or learning algorithms. Instead, how the learning problem is formulated and how the ML model is applied within the system are equally important factors in achieving practical performance gains.
@@ 7 Conclusion
Acknowledgements: We want to thank Larry Greenfield, Tzu-Wei Yang, Deniz Altınbüken and the anonymous reviewers for invaluable feedback and discussions.
@@ C.2 Sensitivity Analysis
Adaptive Algorithm Parameters. We include all combinations of hyperparameters where $T_{\textsc{SpilloverTCIO}}\in\{[0.005,$ 0.03],0.03],0.03 ] , [0.01,[0.01,[ 0.01 , 0.15],[0.05,0.25]}0.15],[0.05,0.25]\}0.15 ] , [ 0.05 , 0.25 ] }, look back window time length (seconds) $t_{w}\in\{600,900,1800\}$, and admission decision effective time $t_{l}\in\{600,900,$ 1800}1800\}1800 }. We evaluate the sensitivity of the TCO savings for the same set of workloads in fig. 8. For each parameter combination, we apply the same parameter settings to all the workloads in the group. In fig. 15, the blue area in the figure presents the upper bound and lower bound of TCO savings under different SSD capacities across different hyperparameter combinations. Our solution is not sensitive in terms of hyperparameter selection in the adaptive algorithm.
@@ C.3 Adaptive Category Selection Dynamics
To demonstrate the dynamics of our adaptive algorithm, we present the pattern of category threshold change and spill over percentage in fig. 16. We track the threshold change for 1 week online. Our adaptive category selection algorithm can adjust the category admission threshold to a higher range when SSD quota is limited and allow more category admissions when SSD space is plentiful.
@@ Appendix E Discussion
Our “bring your own model” idea can be adopted in other deployments and frameworks. The solution flow of our system is: 1) Setting an optimization objective (TCO savings in our case). 2) Designing a ‘hint’ (workload model output) passed from each workload from the application layer to the caching layer. Our hint is job importance in terms of TCO savings. 3) Picking jobs of different categories based on feedback from system utilization.
This BYOM flow remains adaptable to other applications. The key components that vary are the objective function, available model features, and how system utilization is quantified.